# M3-20M: A Large-Scale Multi-Modal Molecule Dataset for AI-driven Drug Design and Discovery

### v1

@@ Abstract
This paper introduces M3-20M, a large-scale Multi-Modal Molecular dataset that contains over 20 million molecules. Designed to support AI-driven drug design and discovery, M3-20M is 71 times more in the number of molecules than the largest existing dataset, providing an unprecedented scale that can highly benefit training or fine-tuning large (language) models with superior performance for drug design and discovery. This dataset integrates one-dimensional SMILES, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions collected through web crawling and generated by using GPT-3.5, offering a comprehensive view of each molecule. To demonstrate the power of M3-20M in drug design and discovery, we conduct extensive experiments on two key tasks: molecule generation and molecular property prediction, using large language models including GLM4, GPT-3.5, and GPT-4. Our experimental results show that M3-20M can significantly boost model performance in both tasks. Specifically, it enables the models to generate more diverse and valid molecular structures and achieve higher property prediction accuracy than the existing single-modal datasets, which validates the value and potential of M3-20M in supporting AI-driven drug design and discovery. The dataset is available at https://github.com/bz99bz/M-3.
@@
American Chemical Society,
[…]
{tocentry}
@@ 1 Introduction
Drug design and discovery is an important field in pharmacy, which aims to identify new therapeutic compounds and optimize their properties for clinical use 1, 2, 3, 4, 5. This process involves the prediction of molecule-target interactions 6, 7, 8, 9 and molecular properties 10, 11, 12, 13, and the design or generation of targeted molecules 14, 15, 16, 17, 18, which are fundamental for developing effective and safe drugs. In recent years, various models trained or pre-trained and fine-tuned using massive data, such as 19, 20, 21, 22, 23, 24, 25, 26, 27, have emerged as powerful tools in this domain. These models leverage large datasets and sophisticated algorithms to generate new compounds, predict molecular structures, and analyze their biochemical properties, significantly advancing the process of drug design and discovery. However, most existing molecular datasets for model training are limited to single-modality, failing to capture the characteristics of molecules completely, thus hampering the building of powerful models for drug design and discovery.
@@
Recent works 23, 28, 29 have been paying increasing attention to multi-modal molecule datasets. Zeng 29 introduced a deep-learning system that bridges molecular structures and biomedical texts, achieving comprehension comparable to human professionals. Liu et al. 28 developed a multi-modal model for text-based retrieval and editing of molecular structures, facilitating advanced molecular text analysis. Liu et al. 23 presented a multi-modal large language model for molecular science that integrates graphs, images, and texts to support various downstream tasks. However, datasets used in these works still suffer from significant limitations. On the one hand, these datasets are typically small in size, encompassing a limited chemical space that restricts the generalization power of the trained or tuned models. On the other hand, the absence of complete modalities of molecular data undermines the performance of computational models.
@@
To address these limitations, this paper presents a new and large-scale multi-modal integrated molecule dataset called M3-20M for AI-driven drug design and discovery. M3-20M contains more than 20 million molecules with their SMILES strings, 2D graphs, 3D structures, physicochemical properties, and textual descriptions, providing unprecedented data scale, diversity, and comprehensiveness for training models of superior performance for various drug design and discovery downstream tasks. Concretely, M3-20M is 71 times more in the number of molecules than the largest existing dataset28. It collects physicochemical properties from the PubChem database 30 and enriches the description texts of molecules by both web crawling and GPT-3.5 generation. Therefore, each molecule’s SMILES corresponds to its 2D graph and 3D structure, physicochemical properties, and description texts. Figure 1 shows some examples from our M3-20M dataset.
@@
In summary, M3-20M stands out of the existing datasets in at least the following three aspects: 1) Large scale: it contains over 20 million molecules, which is the largest open-access multi-modal molecule dataset for AI-driven drug design and discovery, to the best of our knowledge. 2) Comprehensive modalities: M3-20M boasts a more complete range of modalities, including one-dimensional molecular SMILES strings, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions. Such multi-modal data offers a holistic perspective of each molecule, benefiting model training and tuning for drug design and discovery. Additionally, we offer tools to generate 2D molecular graph images that facilitate compound patent retrieval, and to crawl PubMed for research articles related to specific molecules, thereby enriching textual descriptions. 3) Supporting various tasks: M3-20M can support model training and tuning for various downstream drug design and discovery tasks, including molecule generation, molecular property prediction, lead optimization, and virtual screening. It also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction. Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
@@
To demonstrate the power of M3-20M in supporting drug design and discovery, extensive experiments on two basic tasks, molecule generation and molecular property prediction, are conducted using large language models including GLM4 31, GPT-3.5 32, and GPT-4 33. The experimental results show that M3-20M can significantly boost model performance in both tasks. Concretely, with M3-20M, the above models can generate more diverse and valid molecular structures, and achieve higher property prediction accuracy than with existing single-modal datasets, thus verifying the value and potential of M3-20M in supporting AI-driven drug design and discovery.
@@ 2 Related Work
Multi-modal molecular data is becoming increasingly important in artificial intelligence and molecular biology. which can provide more comprehensive and rich information, bringing new opportunities and challenges to drug discovery, molecular property prediction, and generation tasks. For example, Wang et al. 34 pointed out that the use of multi-modal data, including sequence, structure, text etc., can improve the accuracy of drug discovery. However, the challenges faced by multi-modal data cannot be ignored, as different data types often necessitate distinct methods for organization, storage and processing as well as learning. Additionally, ensuring data consistency and integrity also poses a critical issue, with various forms of data potentially exhibiting semantic gaps or inconsistencies. Therefore, there is an urgent need to overcome the above issues to create a large-scale multi-modal integrated dataset for model training or tuning in drug design and discovery. In what follows, we briefly review the related work in three perspectives, molecular representation, single and two-modality molecule datasets, and multi-modality molecule datasets.
@@ 2.1 Molecular representations
Molecules can be represented by various modalities of information 35, 36, including one-dimensional sequences, two-dimensional graphs or images, three-dimensional structures, and molecular properties as well as textual descriptions. One-dimensional sequences primarily consist of SMILES strings, molecular fingerprints, and algebraic topology-based characterizations. Two-dimensional representations include molecular graphs and images generated using toolkits such as RDKit 37 or Open Babel 38. Three-dimensional structures are categorized into molecular graphs and grids 39, containing atomic coordinates that provide detailed spatial information. Molecular textual descriptions offer additional information derived from experiments. Single-modal data is a reduction of the complete structural information of the molecules, which means the representations learned from single-modal suffer from information loss. For example, two SMILES may represent the same molecule or a two-dimensional molecular graph cannot distinguish between the conformations. Therefore, integrating multi-modal information is crucial for molecular modeling and model training in drug design and discovery.
@@ 2.2 Single-modal and two-modal molecule datasets
Most existing molecule datasets are limited to single or dual modalities, restricting the generalization power of trained models 40, 41, 42, 43, 44. These datasets typically have only one-dimensional (1D) SMILES strings, or two-dimensional (2D) structures, or three-dimensional (3D) structures, and usually lack textual molecular descriptions. For instance, MOSES 45 comprises 4,591,276 molecules filtered by molecular weight, ranging from 250 to 350 Daltons, and contains solely one-dimensional SMILES representations. Similarly, ZINC 46, which serves as a repository of commercially available compounds for virtual screening, contains over 1.3 billion molecules obtained from 310 catalogs across 150 vendors, encompassing 2D graphs and, for the majority, 3D structures. Zeng et al. 29 presented a deep learning system that connects molecular structures with biomedical texts, demonstrating a level of comprehension comparable to that of human experts.
@@ 2.3 Multi-modal molecule datasets
With the development of AI-driven drug design and discovery, multi-modal data have become increasingly important in recent years. Up to now, only a few datasets have been introduced to address this need, each with distinct purposes and contributions.
Liu et al. 28 constructed PubChemSTM, a dataset containing over 280,000 chemical structure–text pairs. Despite PubChemSTM’s ability to generalize novel biochemical concepts across various benchmarks, its dataset volume remains insufficient. Subsequently, the igcdata dataset 23 was proposed, which comprises 220,000 descriptions sourced from PubChem and ChEBI-20 47. Although these two datasets provide valuable resources, their relatively small scale restricts their utility for training high-performance models. Beyond scale, these two datasets also lack the diversity and completeness of molecular modalities necessary for robust model development. Issues such as sparse or overly brief textual descriptions further exacerbate this limitation, leading to suboptimal training outcomes of drug design and discovery models.
To address the limitations of existing datasets and support the training or tuning of high-performance models for drug design and discovery, we present the M3-20M dataset, which significantly expands the scale and scope of multi-modal molecular data. Table 1 offers a qualitative comparison of our dataset against existing multi-modal datasets across seven dimensions, including the number of molecules contained, whether consisting of 3D structures, 2D graphs, PubChem CID, physicochemical properties and downstream task datasets, and the number of description texts. From Table 1, we can see that our dataset M3-20M has the largest number of molcuels and covers the most comprehensive molecular data modalities.
Property QM948 PCdes29 PubChemSTM28 igcdata23 M3-20M (Ours)
@@
3 Dataset Construction
This section introduces the dataset construction process and some statistic information of the dataset. Molecular data is collected from existing databases PubChem, ZINC and QM9, and then processed to get multiple representations of molecules, including SMILES strings, 2D and 3D structures, physicochemical properties, and textual descriptions. Notably, we enhance molecular textual descriptions by generating using GPT-3.5 with an expert scoring mechanism to ensure that these descriptions are both scientifically accurate and practically useful. All the data are integrated into a comprehensive dataset with over 20 million molecules.
Figure 2: The construction pipeline of our multi-modal molecular dataset M3-20M.
@@ 3.1 Data Collection
Figure 2 shows the pipeline for constructing our multi-modal molecular dataset M3-20M. To obtain molecules and the corresponding molecular graphs and 3D structures, we first collect molecules from the PubChem, ZINC, and QM9 databases for their extensive molecular coverage and reliable data quality. Then, we utilize the Chem function of RDKit to extract atomic features and chemical bond characteristics and convert them into 2D molecular graphs. Additionally, we employ the PubChem API to batch download SDF files containing the 3D structures of molecules and use the GetAtomPosition function of RDKit to calculate the 3D coordinates of the central atoms.
Textual descriptions serve as a bridge for large language models to comprehend multi-modal molecules. The remarkable recognition accuracy and robustness of the CLIP 49 model exemplify the benefits of effectively linking textual descriptions with other modality information of molecules. Similarly, since textual descriptions of molecules facilitate a deeper understanding of chemical information, we meticulously construct these descriptions in the dataset using three approaches: 1) For the molecules that have textual descriptions in PubChem, we directly extract the texts from PubChem. 2) For the molecules that do not have textual descriptions but have physicochemical properties in PubChem, we transform these physicochemical property values into textual descriptions, which will be detailed later. 3) For those without physicochemical property values or not present in the PubChem database, we generate their textual descriptions by GPT-3.5. We present the details in the following sections.
Initially, we utilized textual data from PubChem; however, only a subset of molecules in PubChem contains associated textual descriptions. The sparsity of text information leads to poor performance of the trained or tuned models. So we resort to text augmentation. To generate textual descriptions for molecules having physicochemical property values in PubChem, we extract 26 key physicochemical properties and their values from PubChem, including Molecular Weight, XLogP3, Hydrogen Bond Donor Count, Hydrogen Bond Acceptor Count, Rotatable Bond Count, Exact Mass, Monoisotopic Mass, Topological Polar Surface Area, Heavy Atom Count, Formal Charge, Complexity, Isotope Atom Count, Defined Atom Stereocenter Count, Undefined Atom Stereocenter Count, Defined Bond Stereocenter Count, Undefined Bond Stereocenter Count, Covalently-Bonded Unit Count, Canonicalization Status, Physical Description, Color/Form, Odor, Boiling Point, Melting Point, Flash Point, and Solubility Density.
@@
Furthermore, we expand seven downstream task datasets with multi-modal information from our dataset by associating the PubChem CIDs of molecules, including molecule generation datasets MOSES and QM9, and molecular property prediction datasets such as BBBP, BACE, HIV, Tox21, and ClinTox. This expansion enables model enhancement with multi-modal information for both molecule generation and property prediction tasks. The expanded datasets are denoted as MOSES-MM, QM9-MM, BBBP-MM, BACE-MM, HIV-MM, Tox21-MM, and ClinTox-MM, respectively.
@@
As mentioned above, for molecules lacking physicochemical properties or not present in PubChem, we generate molecular textual descriptions as a molecule captioning task by using the GPT-3.5 API. Figure 3 illustrates the process of molecular description text generation. Initially, we use the following prompt to guide GPT-3.5 in generating descriptions based solely on SMILES strings: “You are an expert chemist. Given the molecular SMILES, your task is to provide a detailed description of the molecule using your experienced chemical knowledge.”
@@
However, this approach results in issues such as incorrect naming of SMILES strings and misidentifying functional groups. To address these problems, inspired by the CO-STAR (Context-Objective-Style-Tone-Audience-Response) framework, we rewrite the original prompt as follows to enhance the output quality: “You are an expert chemist with a comprehensive understanding of molecular structures and chemistry. Using your extensive knowledge, please describe the molecule represented by the following SMILES string. Focus on its structural features, potential chemical properties, and any known applications or reactions it may participate in. Use only the information inferred from the SMILES representation; do not create or infer any hypothetical or novel names for the molecule.”
Utilizing this enhanced prompt, we generated 1,073,845 descriptions using GPT-3.5, which is only 0.934% of the total descriptions. The proportion of synthetic data to real data is very low and will not have a negative impact on real data. The following section elaborates on the quality control measures implemented through the Human Expert Scoring Mechanism to ensure the precision and reliability of the generated descriptions.

### v2

@@ Abstract
This paper introduces M3-20M, a large-scale Multi-Modal Molecule dataset that contains over 20 million molecules, with the data mainly being integrated from existing databases and partially generated by large language models. Designed to support AI-driven drug design and discovery, M3-20M is 71 times more in the number of molecules than the largest existing dataset, providing an unprecedented scale that can highly benefit the training or fine-tuning of models, including large language models for drug design and discovery tasks. This dataset integrates one-dimensional SMILES, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions collected through web crawling and generated using GPT-3.5, offering a comprehensive view of each molecule. To demonstrate the power of M3-20M in drug design and discovery, we conduct extensive experiments on two key tasks: molecule generation and molecular property prediction, using large language models including GLM4, GPT-3.5, GPT-4, and Llama3-8b. Our experimental results show that M3-20M can significantly boost model performance in both tasks. Specifically, it enables the models to generate more diverse and valid molecular structures and achieve higher property prediction accuracy than existing single-modal datasets, which validates the value and potential of M3-20M in supporting AI-driven drug design and discovery. The dataset is available at https://github.com/bz99bz/M-3.
@@
Multi-modal Molecule Dataset, Dataset construction, Molecule generation, Molecule property prediction
@@ 1 Introduction
As an important field in pharmacy, drug design and discovery aim to identify novel therapeutic compounds and optimize their properties for clinical application santos2017comprehensive ; hughes2011principles ; tsaioun2016evidence ; wu2018moleculenet ; chen2023artificial . This process involves a series of tasks, from the prediction of molecule-target interactions ji2023comprehensive ; ye2021unified ; li2022effective ; lu2024dynamicbind and molecular properties fu2024admetlab ; goldman2023prefix ; rong2020self ; masters2022gps++ , to the design or generation of targeted molecules gao2022sample ; franke2022probabilistic ; kong2022molecule ; yang2021hit ; song2024equivariant , which are fundamental for developing effective and safe drugs. In recent years, various models trained or pre-trained and fine-tuned using massive data, such as bagal2021molgpt ; chithrananda2020chemberta ; honda2019smiles ; fabian2020molecular ; liu2024git ; guo2023indeed ; liu2023group ; ai2024extracting ; yu2023multimodal ; ye2025drugassist , have emerged as powerful tools in this domain. These models leverage large datasets and sophisticated algorithms to generate new compounds, predict molecular structures, and analyze their biochemical properties, significantly accelerating drug design and discovery. However, most existing molecule datasets for model training are limited to single-modality, failing to capture molecular characteristics comprehensively, thus hindering the development of powerful models for drug design and discovery
@@
Recent works liu2024git ; liu2023multi ; zeng2022deep are paying increasing attention to multi-modal molecule datasets. Zeng zeng2022deep introduced a deep-learning system that bridges molecular structures and biomedical texts, achieving comprehension capacity comparable to human professionals. Liu et al. liu2023multi developed a multi-modal model for text-based retrieval and editing of molecular structures, facilitating advanced molecular text analysis. Liu et al. liu2024git presented a multi-modal large language model for molecular science that integrates graphs, images, and texts to support various downstream tasks. However, datasets used in these works still suffer from significant limitations. On the one hand, these datasets are typically small in size, encompassing a limited chemical space that restricts the generalization power of the trained or tuned models. On the other hand, the absence of complete modalities of molecular data undermines the performance of trained models.
@@
To address these limitations, this paper presents a new and large-scale multi-modal integrated molecule dataset called M3-20M for AI-driven drug design and discovery. Data in M3-20M are mainly integrated from existing molecular databases, and supplemented by generation with large language models (LLMs). M3-20M contains more than 20 million molecules with their SMILES strings, 2D graphs, 3D structures, physicochemical properties, and textual descriptions, providing unprecedented data scale, diversity, and comprehensiveness for training models of drug design and discovery downstream tasks. M3-20M is 71 times more in the number of molecules than the largest existing multi-modal dataset PubChemSTM liu2023multi . It collects physicochemical properties from the PubChem database kim2016pubchem and enriches the description texts of molecules by web crawling and GPT-3.5 generation. Therefore, each molecule’s SMILES corresponds to its 2D graph and 3D structure, physicochemical properties, and description texts. Figure 1 shows some examples from our M3-20M dataset.
@@
In summary, M3-20M stands out of the existing datasets in at least the following three aspects: 1) Large scale: it contains over 20 million molecules, which is the largest open-access multi-modal molecule dataset for AI-driven drug design and discovery, to the best of our knowledge. 2) Comprehensive modalities: M3-20M boasts a more complete range of modalities, including one-dimensional molecular SMILES strings, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions. Such multi-modal data offers a holistic perspective of each molecule, benefiting model training and tuning for drug design and discovery. Additionally, we offer tools to generate 2D molecular graph images that facilitate compound patent retrieval, and to crawl PubMed for research articles related to specific molecules, thereby enriching textual descriptions. 3) Supporting various tasks: M3-20M can support model training and tuning for various downstream drug design and discovery tasks, including molecule generation, molecular property prediction, lead optimization, and virtual screening. It also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction. Moreover, we construct seven multi-modal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
@@
To demonstrate M3-20M’s value and potential in supporting drug design and discovery, we conduct extensive experiments on two fundamental tasks, molecule generation and molecular property prediction. Our evaluation encompasses three categories of models, including closed-source LLMs (GLM4 du2022glm , GPT-3.5 ouyang2022training , and GPT-4 achiam2023gpt ), open-source LLM (Llama3-8b dubey2024llama ) and specialized molecule generation models (MoFlow zang2020moflow and D2L-OMP guo2024diffusing ). Our experimental results show that M3-20M can significantly boost model performance in both tasks. Concretely, with M3-20M, the above models can generate more diverse and valid molecular structures, and achieve higher property prediction accuracy than using existing single-modal datasets.
@@ 2 Related work
Multi-modal molecular data offers comprehensive AI and molecular biology insights, enhancing drug discovery, property prediction, and generation tasks. As Wang et al. wang2023multitask demonstrated, integrating sequence, structure, and text data improves drug discovery accuracy. However, challenges persist in organization, processing, and maintaining consistency across different modalities. There’s an urgent need for large-scale integrated multi-modal datasets for drug design model development. In what follows, we briefly review the related work in three perspectives, molecular representation, single and two-modality molecule datasets, and multi-modality molecule datasets.
@@ 2.1 Molecular representations
Molecules can be represented through diverse modalities david2020molecular ; ertl2010molecular : 1D sequences (SMILES, fingerprints, topological characterizations), 2D visualizations (graphs, RDKit/Open Babel landrum2013rdkit ; o2011open images), 3D structures (atomic coordinate graphs/grids li2022deep ), and textual descriptions from experiments. Single-modal data is a reduction of the complete structural information of the molecules, which means the representations learned from single-modal suffer from information loss. For example, two SMILES may represent the same molecule or a two-dimensional molecular graph cannot distinguish between the conformations. Therefore, integrating multi-modal information is crucial for molecular modeling and model training in drug design and discovery.
@@ 2.2 Single-modal and two-modal molecule datasets
Existing molecule datasets are predominantly limited to single or dual modalities (1D SMILES strings, 2D structures, or 3D conformations), restricting the generalization power of trained models wang2023multimodal ; fink2007virtual ; wang2022deep ; axelrod2022geom ; yang2024quandb . For instance, MOSES polykovskiy2020molecular comprises 4,591,276 molecules with only SMILES representations, while ZINC irwin2005zinc includes over 1.3B molecules with 2D graphs and some 3D structures. Although recent advances like Zeng et al. zeng2022deep presented a molecular structure-text system. Beaini et al. beaini2023towards developed Graphium, a vast repository of 100M molecules with 13B labels, yet it focuses on multi-task learning rather than multi-modal representation.
@@ 2.3 Multi-modal molecule datasets
The growth of AI-driven drug design and discovery has elevated the importance of multi-modal molecular data. Current multi-modal datasets remain limited: PubChemSTM liu2023multi contains 280,000 structure-text pairs but lacks sufficient volume; igcdata liu2024git includes 220,000 descriptions from PubChem and ChEBI-20 gaulton2012chembl ; and ChEBI-20-MM liu2025quantitative provides 32,998 molecules as an evaluation benchmark rather than a training dataset. These datasets suffer from inadequate scale, limited modality coverage, and sparse textual descriptions.
To overcome these limitations, we present the M3-20M dataset, which significantly expands the scale and scope of multi-modal molecular data. Table 1 compares our dataset with existing multi-modal datasets across seven dimensions: molecule count, 3D structures, 2D graphs, PubChem CID, physicochemical properties, downstream tasks, and description texts. M3-20M has the largest number of molecules and covers the most comprehensive molecular data modalities.
Property QM9ramakrishnan2014quantum PCdeszeng2022deep PubChemSTMliu2023multi igcdataliu2024git M3-20M (Ours)
@@
3 Methodology
This section outlines our dataset construction process and statistics. Molecular data is collected from existing databases PubChem, ZINC, and QM9ruddigkeit2012enumeration ; ramakrishnan2014quantum , processing them into multiple representations of molecules, including SMILES strings, 2D and 3D structures, physicochemical properties, and textual descriptions. Notably, we enhance molecular descriptions using GPT-3.5 with an expert scoring mechanism to ensure scientific accuracy. The final dataset integrates over 20 million molecules.
Figure 2: The construction pipeline of our multi-modal molecule dataset M3-20M.
@@ 3.1 Data Collection
Figure 2 illustrates our M3-20M construction pipeline. To obtain molecules and the corresponding molecular graphs and 3D structures, we first collect molecules from the PubChem, ZINC, and QM9 databases for their extensive molecular coverage and reliable data quality. Then, we utilize the Chem function of RDKit to extract atomic features and chemical bond characteristics and convert them into 2D molecular graphs. Additionally, we employ the PubChem API to batch download SDF files containing the 3D structures of molecules and use the GetAtomPosition function of RDKit to calculate the 3D coordinates of the central atoms.
Textual descriptions serve as a bridge for large language models to comprehend multi-modal molecules. For textual descriptions, we employ three approaches: 1) For the molecules that have textual descriptions in PubChem, we directly extract the texts from PubChem. 2) For the molecules that do not have textual descriptions but have physicochemical properties in PubChem, we transform these physicochemical property values into textual descriptions, which will be detailed later. 3) For those without physicochemical property values or not present in the PubChem database, we generate their textual descriptions by GPT-3.5. We present the details in the following sections.
Initially, we utilized textual data from PubChem; however, only a subset of molecules in PubChem contains associated textual descriptions. The sparsity of text information leads to poor performance of the trained or tuned models. So we resort to text augmentation. To generate textual descriptions for molecules having physicochemical property values in PubChem, we extract 26 key physicochemical properties and their values from PubChem (Table 10).
@@
Additionally, we expand seven downstream task datasets (MOSES, QM9, BBBP, BACE, HIV, Tox21, ClinTox) with multi-modal information and the PubChem CIDs, creating their ”-MM” counterparts to enhance both molecule generation and property prediction tasks.
@@
For molecules lacking physicochemical properties or not present in PubChem, we generate textual descriptions using GPT-3.5. Figure 3 illustrates the process of molecular description text generation. Initially, we used a basic prompt identifying GPT-3.5 as an expert chemist, but this produced errors in SMILES interpretation and functional group identification.
@@
Inspired by the CO-STAR (Context-Objective-Style-Tone-Audience-Response) framework, we develop an enhanced prompt: “You are an expert chemist with a comprehensive understanding of molecular structures and chemistry. Using your extensive knowledge, please describe the molecule represented by the following SMILES string. Focus on its structural features, potential chemical properties, and any known applications or reactions it may participate in. Use only the information inferred from the SMILES representation; do not create or infer any hypothetical or novel names for the molecule.”
Utilizing this enhanced prompt, we generated 1,073,845 descriptions, which is only 0.934% of the total descriptions. The proportion of synthetic data to real data is very low and will not have a negative impact on real data. The subsequent section details our Human Expert Scoring Mechanism for quality control.